OpenStudy (zistoe):

The following formula relates the quantities temperature in Celsius (C) and temperature in kelvins (K). C = K − 273.15 Write two other related equations. If the temperature is 23 in degrees Celsius, what is the temperature in kelvins? Show your work. Which is greater, 300 kelvins or 25 degrees Celsius? Show your work. Answer:

OpenStudy (kevin):

A = B - C it same with A + C = B B - A = C

OpenStudy (studygurl14):

1: rearrange the variables by adding/subtracting 2: plug in 23 for C and solve for K 3: Plug in 300 for K and solve for C. Compare to 25 C

OpenStudy (zistoe):

can u help me solve the question step by step

OpenStudy (kevin):

Do you understand what I wrote above?

OpenStudy (zistoe):

no

OpenStudy (kevin):

C = K − 273.15 It's same with C + 273.15 = K - 237.15 + 237.15 --> do you understand this?

OpenStudy (kevin):

adding both of sides using 273.15

OpenStudy (zistoe):

ok whats next

OpenStudy (zistoe):

because i need the answer step by step for the problem who ever answers get the gold medal